40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Message ID: 148     Entry time: Sat Jul 6 17:10:07 2013
Author: Koji 
Type: Mechanics 
Category: Characterization 
Subject: PZT Response analysis 

Analysis of the PZT scan / TF data taken on May 31st and Jun 1st.

[DC scan]

Each PZT was shaken with 10Vpp 1Hz triangular voltage to the thorlabs amp.
The amp gain was x15. Abut 4 TEM00 peaks were seen on a sweep between 0 and 10V.

The input voltage where the peaks were seen was marked. Each peak was mapped on the
corresponding fringe among four. Then the each slope (up and down) was fitted by a iiner slope.
Of course, the PZTs show hystersis. Therefore the result is only an approximation.

PZT1: PZT #26, Mirror C6 (CM1)
PZT2: PZT #23, Mirror C5 (CM2)

PZT arrangement [ELOG Entry]

PZT1:
Ramp Up        13.21nm/V
Ramp Down   13.25nm/V
Ramp Up        13.23nm/V
Ramp Down   13.29nm/V

=> 13.24+/-0.02 nm/V

PZT2:
Ramp Up        13.27nm/V
Ramp Down   12.94nm/V
Ramp Up        12.67nm/V
Ramp Down   12.82nm/V

=> 12.9+/-0.1 nm/V

[AC scan]

The OMC cavity was locked with the fast laser actuation. Each PZT was shaken with a FFT analyzer for transfer function measurments.
(No bias voltage was given)

The displacement data was readout from the laser fast feedback. Since the UGF of the control was above 30kHz, the data was
valid at least up to 30kHz. The over all calibration of the each curve was adjusted so that it agrees with the DC response of the PZTs (as shown above).

The response is pretty similar for these two PZTs. The first series resonance is seen at 10kHz. It is fairly high Q (~30).

Attachment 1: PZT_Scan.pdf  20 kB  Uploaded Sat Jul 6 18:35:15 2013  | Hide | Hide all | Show all
PZT_Scan.pdf
Attachment 2: L1OMC_PZT_Response.pdf  47 kB  Uploaded Sat Jul 6 18:38:57 2013  | Show | Hide all | Show all
ELOG V3.1.3-